OpenStudy (anonymous):

Solve By Substitution..... { 1/3x - 2/3 y-1=0 { x+y+7=0

OpenStudy (anonymous):

in the 1st equation, u will get 1/3x=2/3y+1 or, x=2y+3 in the second equation, substitute x with 2y+3 then u will get y=-10/3 put the value of y in any equation, u will get the solution for x

OpenStudy (anonymous):

{ 1/3x - 2/3 y-1=0 (1) { x+y+7=0 (2) ------> x=-y-7 subs this to eq 1 (1/3)x -( 2/3)y -1=0 (1/3)(-y-7) -( 2/3)y -1=0 mult by 3 (-y-7) - 2y -3=0 -3y-10=0 -3y=10 y=-10/3 ans sub this to eq (2) x+y+7=0 x-(10/3)+7=0 x=-7 +10/3 x=-21/3 +10/3 x=-11/3 ans